A per invests 10000 for two years at a certain rate of interest compounded annually. At the end of one year this sum amounts to 11200. Calculate:
(i). the rate of interest per annum
(ii). the amount at the end of second year.

Answer by mananth(16946)

i= (fv/pv)^(1/n) -1

fv= 11200
pv= 10000
n= 1 1.67
i=( 11200 / 10000 )^ 1 / n - 1
i= 1.12 ^ 1 / 1 - 1 9.71
i= 0.03 1.94
i= 3 %
------
Principal P = 10000
Amount= A
years=n 2.00
compounded 1 times a year t
Rate = 3.00 0.03
Amount = P*((n+r)/n)^n*t

Amount = = 10000 *( 1 + 0.03 )^ 2 * 1
Amount = 10000 *( 1 + 0.03 )^ 2
10000 *( 1.03 )^ 2
Amount = $10609